# Complexity of Linear Subsequences of $k$-Automatic Sequences

### v2

@@ 4 Linear Subsequences
We now consider an interesting application of this connection between state complexity and subword complexity. Given an automatic sequence ${\bf h}=(h(i))_{i\geq 0}$ we would like to obtain an upper bound on the subword complexity of its linear subsequence $(h(ni))_{i\geq 0}$. This can be done using our results.
@@
Clearly $\rho_{{\bf h}_{n}}(\ell)\leq\rho_{{\bf h}^{\prime}_{n}}(\ell)$. Let $M=(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$ be an automaton generating ${\bf h}^{\prime}_{n}$ built by the construction in Theorem 10, and let $\hat{\bf M}$ be the sequence generated by the intrinsic DFAO $\hat{M}=(Q,\Sigma_{k},\delta,q_{0},Q,i)$ where $i:Q\rightarrow Q$ is the identity map. By Theorem 2, we have $\rho_{\bf h_{n}^{\prime}}(\ell)\leq k\ell\rho_{\hat{\bf M}}(2)$. However, each length-$2$ subword of $\hat{\bf M}$ corresponds, from the construction of Theorem 10, to a distinct length-$2n$ subword of $\bf h$$^{\prime}$. Thus $\rho_{\hat{\bf M}}(2)\leq\rho_{{\bf h}^{\prime}}(2n)$. Putting this all together gives
@@
A superficially similar notion of complexity was studied by Konieczny and Müllner [16]. However, the particular notion of complexity they studied is the number of length-$l$ subwords of all of the subsequences $(h(ni+c))_{i\geq 0}$ for integers $n\geq 1$ and $c\geq 0$; this notion was first introduced by Avgustinovich et al. [3]. In contrast, the complexity we discuss in this paper is the number of length-$l$ subwords of a single subsequence $(h(ni))_{i\geq 0}$ with fixed $n$.
4.1 The Thue-Morse Sequence
@@
For all integers $n\geq 1$, the number of states in the minimal automaton generating $(t(ni))_{i\geq 0}$ is $\rho_{\bf t}(n/\nu_{2}(n))$.
@@
For an odd integer $n$, all length-$n$ subwords of the Thue-Morse sequence $\bf t$ are reachable states in the automaton generating $(t(ni))_{i\geq 0}$ built by the construction provided in Theorem 10.
@@
We use results by Blanchet-Sadri et al. [5, Lemma 12] by setting $i=0$ in their lemma. As a result, for each subword of the Thue-Morse sequence and odd integer $n$, there exists integer $i^{\prime}$ such that the subword occurs at index $ni^{\prime}$ of the Thue-Morse sequence; in other words, the state corresponding to the subword is reachable by input $(i^{\prime})_{2}$. ∎
@@
For any word $z$, since $t(n[x])=t(n[y])$, we have $t(n[x0^{|z|}])=t(n[y0^{|z|}])$. So if there is word $z$ such that $\left\lfloor\frac{n[z]}{2^{|z|}}\right\rfloor=j$, the number of ones in $(n[x0^{|z|}]+n[z])_{2}$ is the sum of the number of ones in $(n[x]+j)_{2}$ and $(n[z]\bmod 2^{|z|})_{2}$; a similar result holds for $y$ instead of $x$. Since $t(n[x]+j)\neq t(n[y]+j)$, we have $t(n[xz])\neq t(n[yz])$. By iterating different $z$, we get $0\leq\lfloor\frac{n[z]}{2^{|z|}}\rfloor<n$. Therefore, such $z$ always exists. ∎
@@
Therefore, in order to obtain the number of states in the minimal automaton generating $(t(ni))_{i\geq 0}$, we remove powers of $2$ from $n$ to get an odd integer $n^{\prime}=n/\nu_{2}(n)$. The automaton generating $(t(n^{\prime}i))_{i\geq 0}$ is the same as the automaton generating $(t(ni))_{i\geq 0}$ and the number of states in the minimal automaton generating $(t(n^{\prime}i))_{i\geq 0}$ is $\rho_{\bf t}(n^{\prime})$. This proves the theorem. ∎
@@
$\operatorname{sc}(t(2ri))$ $=\operatorname{sc}(t(ri)),$
$\operatorname{sc}(t((2^{p}+r)i))$ $=3\cdot 2^{p}+4(r-1),$ $\text{if $1<r<2^{p-1}$ and $r\bmod 2=1$},$
$\operatorname{sc}(t((2^{p}+2^{p-1}+r)i))$ $=5\cdot 2^{p}+2(r-1),$ $\text{if $1<r<2^{p-1}$ and $r\bmod 2=1$}.$
@@
The subword complexity of the Thue-Morse sequence has been studied by Brlek [7], de Luca and Varricchio [17], and Avgustinovich [4]. Brlek [7] proved that for $n\geq 3$ where $n=2^{p}+r^{\prime}+1$ and $0<r^{\prime}\leq 2^{p}$ we have
$\rho_{\bf t}(n)=\begin{cases}6\cdot 2^{p-1}+4r^{\prime},&\text{if $0<r^{\prime}\leq 2^{p-1}$},\\ 8\cdot 2^{p-1}+2r^{\prime},&\text{if $2^{p-1}<r^{\prime}\leq 2^{p}$}.\end{cases}$
By Theorem 15, from the $0<r^{\prime}\leq 2^{p-1}$ case we get the $\operatorname{sc}(t((2^{p}+r)i))=3\cdot 2^{p}+4(r-1)$ and from the $2^{p-1}<r^{\prime}\leq 2^{p}$ case we get the $\operatorname{sc}(t((2^{p}+2^{p-1}+r)i))=5\cdot 2^{p}+2(r-1)$. ∎
@@
At first glance it might appear that Charlier et al. [8, 9] have already proved results similar to those in Corollary 18. However, this is not the case. The DFAO we considered takes $(i)_{2}$ as input and outputs $t(ni)$. The DFA they considered takes $x$ as input and outputs whether $x]_{k}=ni+r$ where $t(i)=0$, $k$ is a power of $2$, and $0\leq r<n$.
We now turn to shifts of the Thue-Morse sequence. We apply Theorem 10 to the Thue-Morse sequence $(t(i))_{i\geq 0}$ and $n=1$.
@@
For all integers $c\geq 1$, the number of states in the minimal automaton generating $(t(i+c))_{i\geq 0}$ is at most $\frac{10}{3}c$.
@@
Unlike the case of Theorem 17, where the construction in Theorem 10 gives the minimal automaton, in the case of Theorem 20, the construction in Theorem 10 does not necessarily give the minimal automaton, so the upper bound in Theorem 20 is not tight. The next theorem gives a lower bound on the state complexity of $(t(i+c))_{i\geq 0}$.
@@
The minimal msd-first DFAO computing $(t(i+c))_{i\geq 0}$ has more than $c^{0.694}$ states for infinitely many $c$.
@@
The idea is to consider the case $c=2^{n}$ for $n\geq 1$. In this case we show, using the Myhill-Nerode theorem, that the minimal DFAO for $(t(i+c))_{i\geq 0}$ has at least $F_{n+1}$ states. We let $S_{n}$ be the set of all binary words of length $n$ that start with 0 and have no two consecutive $1$’s. As is well-known, this set has cardinality $F_{n+1}$. We now show the members of this set are pairwise inequivalent under Myhill-Nerode.
@@
Finally, from the estimate $F_{n+1}\geq{1\over 2}\alpha^{n}$ (which is easily proved by induction), where $\alpha=(1+\sqrt{5})/2\doteq 1.61803$ is the golden ratio, we see that if $c=2^{n}$ then the minimal DFAO has at least
@@
Problem 1.
What is a good formula for the exact number of states in the minimal automaton generating $(t(i+c))_{i\geq 0}$, as a function of $c$?
@@ 5 Construction by Büchi Arithmetic
We now re-examine the previous relations and operations in light of their implementation in an interpretation of Büchi arithmetic (as, for example, in the system Walnut [18]) to find the time required for computing their automata. We use the number of automaton transitions traversed during the computation as a measure of the time required.
@@
Furthermore, the minimization steps in the implementation for an $n$-state automaton can use an algorithm running in $O(n\log n)$ time by Hopcroft [13] or Valmari [19]. The algorithm by Hopcroft can be slightly modified to be applied to DFAOs as described by van Spaendonck [20]. Note that in an $n$-state DFA, the number of transitions is $O(n)$. We assume that the minimization is applied to each DFA or DFAO created.

### v3

@@ 4 Linear Subsequences of Automatic Sequences
We do not know if the bounds in Corollary 11 are tight. This seems worthy of further study, so we state it as an open problem.
Open Problem 1.
Find a class of automatic sequences $(h_{m}(i))_{i\geq 0}$, generated by a DFAO of $m$ states with msd-first input, such that there is a constant $C$ so the linear subsequences $(h_{m}(ni))_{i\geq 0}$ have at least the state complexity $Cnm^{2}$ for infinitely many $n$.
4.1 Connection with Subword Complexity
We now consider an interesting application of this connection between state complexity and subword complexity. Given an automatic sequence ${\bf h}=(h(i))_{i\geq 0}$ we would like to obtain an upper bound on the subword complexity of its linear subsequence $(h(ni))_{i\geq 0}$.
@@ 4.1 Connection with Subword Complexity
Clearly $\rho_{{\bf h}_{n}}(\ell)\leq\rho_{{\bf h}^{\prime}_{n}}(\ell)$. Let $M=(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$ be an automaton generating ${\bf h}^{\prime}_{n}$ built by the construction in Theorem 10, and let $\hat{\bf M}$ be the sequence generated by the intrinsic DFAO $\hat{M}=(Q,\Sigma_{k},\delta,q_{0},Q,\iota)$ where $\iota:Q\rightarrow Q$ is the identity map. By Theorem 2, we have $\rho_{\bf h_{n}^{\prime}}(\ell)\leq k\ell\rho_{\hat{\bf M}}(2)$. However, each length-$2$ subword of $\hat{\bf M}$ corresponds, from the construction of Theorem 10, to a distinct length-$2n$ subword of $\bf h$$^{\prime}$. Thus $\rho_{\hat{\bf M}}(2)\leq\rho_{{\bf h}^{\prime}}(2n)$. Putting this all together gives
@@
A superficially similar notion of complexity was studied by Konieczny and Müllner [20]. However, the particular notion of complexity they studied is the number of length-$\ell$ subwords of all of the subsequences $(h(ni+c))_{i\geq 0}$ for integers $n\geq 1$ and $c\geq 0$; this notion was first introduced by Avgustinovich et al. [5]. In contrast, the complexity we discuss in this paper is the number of length-$\ell$ subwords of a single subsequence $(h(ni))_{i\geq 0}$ with fixed $n$.
4.2 Thue-Morse Sequence
@@
For all integers $n\geq 1$, the number of states in the minimal DFAO generating $(t(ni))_{i\geq 0}$ with msd-first input is $\rho_{\bf t}(n/\nu_{2}(n))$.
@@
For an odd integer $n$, all length-$n$ subwords of the Thue-Morse sequence $\bf t$ are reachable states in the automaton generating $(t(ni))_{i\geq 0}$ with msd-first input built by the construction provided in Theorem 10.
@@
We use results by Blanchet-Sadri et al. [6, Lemma 12] by setting $i=0$ in their lemma. As a result, for each subword of the Thue-Morse sequence and odd integer $n$, there exists integer $i^{\prime}$ such that the subword occurs at index $ni^{\prime}$ of the Thue-Morse sequence; in other words, the state corresponding to the subword is reachable by input $(i^{\prime})_{2}$. ∎
@@
For any word $z$, since $t(n[x])=t(n[y])$, we have $t(n[x0^{|z|}])=t(n[y0^{|z|}])$. So, if there is word $z$ such that $\left\lfloor n[z]/2^{|z|}\right\rfloor=j$, the number of ones in $(n[x0^{|z|}]+n[z])_{2}$ is the sum of the number of ones in $(n[x]+j)_{2}$ and $(n[z]\bmod 2^{|z|})_{2}$; a similar result holds for $y$ instead of $x$. Since $t(n[x]+j)\neq t(n[y]+j)$, we have $t(n[xz])\neq t(n[yz])$. By iterating different $z$, we get $0\leq\lfloor n[z]/2^{|z|}\rfloor<n$. Therefore, such $z$ always exists. ∎
@@
Therefore, in order to obtain the number of states in the minimal automaton generating $(t(ni))_{i\geq 0}$, we remove powers of $2$ from $n$ to get an odd integer $n^{\prime}=n/\nu_{2}(n)$. The automaton generating $(t(n^{\prime}i))_{i\geq 0}$ is the same as the automaton generating $(t(ni))_{i\geq 0}$ and the number of states in the minimal automaton generating $(t(n^{\prime}i))_{i\geq 0}$ is $\rho_{\bf t}(n^{\prime})=\rho_{\bf t}(n/\nu_{2}(n))$. ∎
@@
$\begin{cases}\operatorname{sc}(t(2ri))=\operatorname{sc}(t(ri)),\\ \operatorname{sc}(t((2^{p}+r)i))=3\cdot 2^{p}+4(r-1),&\text{if $1<r<2^{p-1}$ and $r\bmod 2=1$};\\ \operatorname{sc}(t((2^{p}+2^{p-1}+r)i))=5\cdot 2^{p}+2(r-1),&\text{if $1<r<2^{p-1}$ and $r\bmod 2=1$}.\end{cases}$
@@
The subword complexity of the Thue-Morse sequence was studied by Brlek [8], de Luca and Varricchio [21], and Avgustinovich [4]. Brlek [8] proved that for $n\geq 3$ where $n=2^{p}+r^{\prime}+1$ and $0<r^{\prime}\leq 2^{p}$ we have
$\rho_{\bf t}(n)=\begin{cases}6\cdot 2^{p-1}+4r^{\prime},&\text{if $0<r^{\prime}\leq 2^{p-1}$};\\ 8\cdot 2^{p-1}+2r^{\prime},&\text{if $2^{p-1}<r^{\prime}\leq 2^{p}$}.\end{cases}$
By Theorem 15, from the $0<r^{\prime}\leq 2^{p-1}$ case we get $\operatorname{sc}(t((2^{p}+r)i))=3\cdot 2^{p}+4(r-1)$. From the $2^{p-1}<r^{\prime}\leq 2^{p}$ case we get $\operatorname{sc}(t((2^{p}+2^{p-1}+r)i))=5\cdot 2^{p}+2(r-1)$. ∎
@@
At first glance it might appear that Charlier et al. [10, 11] already proved results similar to those in Corollary 18. However, this is not the case. The DFAO we considered takes $(i)_{2}$ as input and outputs $t(ni)$. The DFA they considered takes $x$ as input and decides whether $x]_{k}=ni+r$ where $t(i)=0$, $k$ is a power of $2$, and $0\leq r<n$.
Next we apply Theorem 10 to the Thue-Morse sequence $(t(i))_{i\geq 0}$ and $n=1$.
@@
For all integers $c\geq 1$, the number of states in the minimal automaton generating $(t(i+c))_{i\geq 0}$ with msd-first input is at most $\frac{10}{3}c$.
@@
Unlike the case of Theorem 17, where the construction in Theorem 10 gives the minimal automaton, in the case of Theorem 20, the construction in Theorem 10 does not necessarily give the minimal automaton, so the upper bound in Theorem 20 is not tight. The next theorem gives a weak lower bound on the state complexity of $(t(i+c))_{i\geq 0}$.
@@
The minimal DFAO computing $(t(i+c))_{i\geq 0}$ with msd-first input has more than $c^{0.694}$ states for infinitely many $c$.
@@
The idea is to consider the case $c=2^{n}$ for $n\geq 1$. In this case we show, using the adaptation of the Myhill-Nerode theorem for DFAOs, that the minimal DFAO for $(t(i+c))_{i\geq 0}$ has at least $F_{n+1}$ states. We let $S_{n}$ be the set of all binary words of length $n$ that start with 0 and have no two consecutive $1$’s. As is well-known, this set has cardinality $F_{n+1}$. We now show the members of this set are pairwise inequivalent under the adaptation of the Myhill-Nerode theorem.
@@
Finally, from the estimate $F_{n+1}\geq{1\over 2}\alpha^{n}$ (which is easily proved by induction), where $\alpha=(1+\sqrt{5})/2\doteq 1.61803$ is the golden ratio, we see that if $c=2^{n}$, then the minimal DFAO has at least
@@
Open Problem 2.
What is a good formula for the exact number of states in the minimal automaton generating $(t(i+c))_{i\geq 0}$ with msd-first input, as a function of $c$?
4.3 Sequences Whose Shifts Have Small State Complexity
A simple counting argument shows that if an automatic sequence $(h(i))_{i\geq 0}$ is not ultimately periodic, then the state complexity of the shifted sequence $(h(i+c))_{i\geq 0}$ cannot be $o((\log c)/(\log\log c))$. We do not know any aperiodic example, automatic or not, that achieves state complexity $O((\log c)/(\log\log c))$ for all $c$-shifts, but we can get almost that close.
Let the infinite word ${\bf p}_{2}=(P_{2}(i))_{i\geq 0}$ be defined as follows.
$P_{2}(i)=\begin{cases}1,&\text{if $i$ is a power of $2$};\\ 0,&\text{otherwise.}\end{cases}$
Theorem 22.
The number of states in the minimal automaton generating $(P_{2}(i+c))_{i\geq 0}$ with base-$2$ msd-first input is $O(\log c)$.
Proof.
We create an automaton $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$ where $\Sigma=\Delta=\{0,1\}$ with $O(\log c)$ states generating ${\bf p}_{2}$. In order to create the automaton, we need to consider two possibilities for $i$ where $P_{2}(i+c)=1$. Let $c^{\prime}$ be $2^{s}-c$ where $2^{s}$ is the smallest power of $2$ greater than $c$.
Case 1: $i+c=2^{j}$ where $|(i)_{2}|\leq|(c)_{2}|$. In this case, the automaton reads $(c^{\prime})_{2}$ in a series of states.
Case 2: $i+c=2^{j}$ where $|(i)_{2}|>|(c)_{2}|$. In this case, the automaton reads one or more $1$s and then reads $(c^{\prime})_{2}$ padded with leading zeros so the length is the same as $(c)_{2}$.
Now we provide a construction for $M$. First, define $\delta(q_{0},0)=q_{0}$. Next, we address Case 1. Let $(c^{\prime})=c^{\prime\prime}_{0}\cdots c^{\prime\prime}_{l}$. For $p\in[0,l]$, let $q^{\prime\prime}_{p}$ be a state in the automaton corresponding to $c^{\prime\prime}_{0}\cdots c^{\prime\prime}_{p}$ read and for $p\in[0,l-1]$, let $\delta(q^{\prime\prime}_{p},c^{\prime\prime}_{p+1})=q^{\prime\prime}_{p+1}$ and $\delta(q_{0},1)=q^{\prime\prime}_{0}$.
Next, we address Case 2. Let $c^{\prime}_{0}\cdots c^{\prime}_{|(c)_{2}|-1}$ be $(c^{\prime})_{2}$ padded with leading 0s so the length is $|(c)_{2}|$. For $p\in[0,|(c)_{2}|-1]$, let $q^{\prime}_{p}$ be a state corresponding to $c^{\prime}_{0}\cdots c^{\prime}_{p}$ read and for $p\in[0,|(c)_{2}|-2]$, let $\delta(q^{\prime}_{p},c^{\prime}_{p+1})=q^{\prime}_{p+1}$. Let $q_{1}$ be the state corresponding to reading the initial $1$s in Case 2. So $\delta(q^{\prime\prime}_{0},1)=q_{1}$, $\delta(q_{1},1)=q_{1}$, and if $c^{\prime}_{0}=1$, then $q_{1}=q^{\prime}_{0}$. Furthermore, we need to add states and transitions so the automaton accepts $1c^{\prime}_{0}\cdots c^{\prime}_{|(c)_{2}|-1}$. Let $p^{\prime}$ be the first index starting from 0 where $c^{\prime\prime}_{0}\cdots c^{\prime\prime}_{l}$ differs from $1c^{\prime}_{0}\cdots c^{\prime}_{|(c)_{2}|-1}$, which can potentially be $l+1$. We set $\delta(q^{\prime\prime}_{p^{\prime}-1},c^{\prime}_{p^{\prime}-1})=q^{\prime}_{p^{\prime}-1}$.
So the DFAO $M=(Q,\Sigma_{2},\delta,q_{0},\Delta,\tau)$ can be formally defined as follows.
$Q$ $=\{q_{0},q_{1},q^{\prime\prime}_{0},\ldots,q_{l}^{\prime\prime},q^{\prime}_{0},\ldots,q^{\prime}_{|(c)_{2}-1|}\},$
[…]
$\Delta$ $=\{0,1\},$
$\tau(q)$ $=\begin{cases}1,&\text{if $q=q^{\prime\prime}_{l}$ or $q=q^{\prime}_{|(c)_{2}|-1}$};\\ 0,&\text{otherwise}.\end{cases}$
The correctness of $M$ is explained above. The number of states in $M$ is $2+(l+1)+(c)_{2}=O(\log c)$. ∎
@@ 5 Construction by Büchi Arithmetic
We now re-examine the previous relations and operations in light of their implementation in an interpretation of Büchi arithmetic (as, for example, in the system Walnut [24]) to find the time required for constructing their automata.
Büchi arithmetic is the logical theory of the natural numbers together with addition and the function $V_{k}(n)$, the largest power of $k$ dividing $n$, for some fixed $k\geq 2$. This theory is powerful enough to express finite automata [9]; in what follows we treat DFAO’s computing automatic sequences as more or less a primitive object, without requiring an explicit expression in terms of $\langle\mathbb{N},+,V_{k}\rangle$.
We use the total number of transitions of an automaton $M$ as a proxy for the computational complexity of constructing $M$. Sometimes the complexity bounds in this section include the addition of $O(\log i)$ or $O(i^{j})$ components for some constant $i$; these addition components are not significant compared to multiplication in $O(\log i)$ or $O(i^{j})$ components for some constant $i$.
In particular, in its current version, Walnut does not use the algorithms for creating the automata we discussed in previous sections. Instead, more general techniques, suitable for any addable numeration system, are used. This means that the size of intermediate automata may differ from the minimal automata we constructed, and the running time can similarly be larger.
@@
Furthermore, the minimization steps in the implementation for an $n$-state automaton can use an algorithm running in $O(n\log n)$ time by Hopcroft [16] or Valmari [26]. The algorithm by Hopcroft can be slightly modified to be applied to DFAOs as described by van Spaendonck [25]. Note that in an $n$-state DFA, the number of transitions is $O(n)$. We assume that the minimization is applied to each DFA or DFAO created without mentioning the minimization step every time.
We start with an analysis of how an automaton for a statement like $x]_{k}=c$ is created, where $c$ is a fixed natural number constant. Recall that in Walnut, constants are entered as numbers written in decimal notation (base $10$). There are well-known efficient algorithms for converting a number from base $10$ to base $k$ for an integer $k\geq 2$; see, for example, [19, §4.4], but these are not used by Walnut. Instead, a more general technique, suitable for any numeration system, is used.
Lemma 23.
The minimal automaton for recognizing the relation $x]_{k}=c$ with msd-first input has $\lfloor\log_{k}c\rfloor+2$ states.
Proof.
Such an automaton consists of a linear chain of nodes labeled with the base-$k$ representation of $c$, which has length $\lfloor\log_{k}c\rfloor+1$. ∎